### v1

@@ Appendix D Proofs for Section 4
The fact that $E^{\prime}<E$ means that the leader’s exploration phase takes place entirely during the follower’s exploration phase. Moreover, since the leader’s exploration parameter $E^{\prime}\cdot|\mathcal{B}|$ is divisible by $|\mathcal{B}|$, for every arm $a\in\mathcal{A}$, the follower pulls every arm $b\in\mathcal{B}$ an equal number of times. Given that follower explores evenly between the two arms $b_{1}$ and $b_{2}$, the leader’s expected average reward $\mathbb{E}[\hat{v}^{1}_{a_{1}}]$ from $a_{1}$ during the first $E^{\prime}\cdot|\mathcal{B}|$ rounds is given by $(0.6+0.2)/2=0.4$ and the leader’s expected average reward $\mathbb{E}[\hat{v}^{1}_{a_{2}}]$ average reward from $a_{2}$ is given by $(0.5+0.4)/2=0.45$.
The proofs boils down to analyzing the relationship between the distributions $\hat{v}^{1}_{a_{1}}$ and $v^{1}_{a_{2}}$. Note that we allow $E,E^{\prime}$ to be arbitrary, so we cannot use standard concentration bounds. Instead, we leverage the symmetry of the distribution of the empirical mean $\hat{v}_{1}(a_{1})$ (this follows from the fact that $\hat{v}_{1}(a_{1})-\mathbb{E}[\hat{v}_{1}(a_{1})]$ is distributed as a Gaussian). This means that:
@@
Assume that the clean event $G$ holds. This occurs with probability at least $1-(|\mathcal{A}|\cdot|\mathcal{B}|+|\mathcal{A}|)T^{-3}$ (Lemma D.1), so the clean event not occuring counts negligibly towards regret.
@@
We focus on $t>E_{2}\cdot|\mathcal{B}|\cdot|\mathcal{A}|+E_{1}\cdot|\mathcal{A}|$ for the remainder of the analysis. Our main ingredient is Lemma D.2. Note that $\epsilon^{*}=\Theta\left(\max\left(\frac{\sqrt{\log T}}{\sqrt{E_{1}}},\frac{\sqrt{\log T}}{\sqrt{E_{2}}}\right)\right)=\Theta\left((|\mathcal{A}|\cdot|\mathcal{B}|\cdot(\log T))^{1/3}\cdot T^{-1/3}\right)$ based on the settings of $E_{1}$ and $E_{2}$. The regret of the leader can be bounded as:
@@
where (A) follows from Lemma D.2. The regret of the follower can similarly be bounded as:
@@
where (B) follows from Lemma D.2. This proves the desired result.
D.3 Proof of Theorem 4.5
@@
Lemma D.3.
Assume the notation above. Let $\texttt{ALG}_{2}$ be any algorithm with high-probability instantaneous regret $g$, and let $\texttt{ALG}_{1}=\texttt{ExploreThenUCB}(E)$. Then, the event $G$ occurs with high probability: $\mathbb{P}[G]\geq 1-T^{-3}(|\mathcal{A}|+1)$.
@@
Lemma D.4.
Assume the notation above. Let $\texttt{ALG}_{2}$ be any algorithm with high-probability instantaneous regret $g$, and let $\texttt{ALG}_{1}=\texttt{ExploreThenUCB}(E)$. Suppose that the clean event $G$ holds. Then it holds that:
$\frac{1}{n_{E\cdot|\mathcal{A}|,T}(a^{\prime})-1}\sum_{t\in S(a^{\prime})\setminus\left\{\max(S(a^{\prime}))\right\}}v_{1}(a_{t},b_{t})\geq\max_{a\in\mathcal{A}}\min_{b\in\mathcal{B}_{\epsilon^{*}}(a)}v_{1}(a,b)-\frac{20\sqrt{\log T}}{\sqrt{n_{E\cdot|\mathcal{A}|,T}(a^{\prime})}-1}.$
@@
$\frac{1}{n_{E\cdot|\mathcal{A}|,T}(a^{\prime})-1}\sum_{t\in S(a^{\prime})\setminus\left\{t^{*}\right\}}v_{1}(a_{t},b_{t})$ $\geq_{(1)}\hat{v}_{1,t^{*}}(a^{\prime})-\frac{10\sqrt{\log T}}{\sqrt{n_{E\cdot|\mathcal{A}|,T}(a^{\prime})-1}}$
$\geq v^{\text{UCB}}_{1,t^{*}}(a^{\prime})-\frac{20\sqrt{\log T}}{\sqrt{n_{E\cdot|\mathcal{A}|,T}(a^{\prime})-1}}$
$\geq v^{\text{UCB}}_{1,t^{*}}(\tilde{a})-\frac{20\sqrt{\log T}}{\sqrt{n_{E\cdot|\mathcal{A}|,T}(a^{\prime})-1}}$
$=\hat{v}_{1,t^{*}}(\tilde{a})+\frac{10\sqrt{\log T}}{\sqrt{|S|}}-\frac{20\sqrt{\log T}}{\sqrt{n_{E\cdot|\mathcal{A}|,T}(a^{\prime})-1}}$
$\geq_{(2)}\frac{1}{|S|}\sum_{t\in S}v_{1}(a_{t},b_{t})-\frac{20\sqrt{\log T}}{\sqrt{n_{E\cdot|\mathcal{A}|,T}(a^{\prime})-1}}$
$\geq_{(3)}\max_{a\in\mathcal{A}}\min_{b\in\mathcal{B}_{\epsilon^{*}}(a)}v_{1}(a,b)-\frac{20\sqrt{\log T}}{\sqrt{n_{E\cdot|\mathcal{A}|,T}(a^{\prime})-1}}$
@@
Assume that the clean event $G$ holds. This occurs with probability at least $1-(1+|\mathcal{A}|)T^{-3}$ (Lemma D.3), so the clean event not occurring counts negligibly towards regret.
@@
$\beta_{1}^{\text{tol}}\cdot(T-E\cdot\left|\mathcal{A}\right|)-\sum_{t=E\cdot|\mathcal{A}|+1}^{T}v^{1}_{a_{t},b_{t}}$
@@
$=\sum_{a\in\mathcal{A}}\sum_{t\in T_{a}}\left(\epsilon^{*}+\max_{a\in\mathcal{A}}\min_{b\in\mathcal{B}_{\epsilon^{*}}(a)}v_{1}(a,b)-v_{1}(a_{t},b_{t})\right)$
$\leq|\mathcal{A}|+\sum_{a\in\mathcal{A}}\sum_{t\in n_{E\cdot|\mathcal{A}|,T}(a)\setminus\left\{\max(S(a))\right\}}\left(\epsilon^{*}+\max_{a\in\mathcal{A}}\min_{b\in\mathcal{B}_{\epsilon^{*}}(a)}v_{1}(a,b)-v_{1}(a_{t},b_{t})\right)$
@@
$+\underbrace{\sum_{a\in\mathcal{A}}(n_{E\cdot|\mathcal{A}|,T}(a)-1)\cdot\left(\max_{a\in\mathcal{A}}\min_{b\in\mathcal{B}_{\epsilon^{*}}(a)}v_{1}(a,b)-\frac{1}{n_{E\cdot|\mathcal{A}|,T}(a)-1}\sum_{t\in S(a)\setminus\left\{\max(S(a))\right\}}\left(v_{1}(a_{t},b_{t})\right)\right)}_{(2)}$
@@
$\sum_{a\in\mathcal{A}}(n_{E\cdot|\mathcal{A}|,T}(a)-1)\cdot\left(\max_{a\in\mathcal{A}}\min_{b\in\mathcal{B}_{\epsilon^{*}}(a)}v_{1}(a,b)-\frac{1}{n_{E\cdot|\mathcal{A}|,T}(a)-1}\sum_{t\in S(a)\setminus\left\{\max(S(a))\right\}}\left(v^{1}_{a_{t},b_{t}}\right)\right)$
$\leq\sum_{a\in\mathcal{A}}(n_{E\cdot|\mathcal{A}|,T}(a)-1)\cdot\frac{20\sqrt{\log T}}{\sqrt{n_{E\cdot|\mathcal{A}|,T}(a)}-1}$
$\leq O\left(\sqrt{\log T}\cdot\sum_{a\in\mathcal{A}}\sqrt{n_{E\cdot|\mathcal{A}|,T}(a)-1}\right)$
@@
where the first inequality uses Lemma D.4 and the last inequality uses Jensen’s inequality.
@@
However, by Lemma D.4, we can also lower bound the average reward across all time steps (except for the last time step) where $a^{\prime}$ is pulled in terms of $n_{E\cdot|\mathcal{A}|,T}(a^{\prime})$ as follows:
@@
$n_{E\cdot|\mathcal{A}|,T}(a^{\prime})\leq\Theta\left(\frac{\log T}{(\epsilon^{*})^{2}}\right)=\Theta\left((\log T)^{1/3}T^{2/3}|\mathcal{A}|^{-2/3}|\mathcal{B}|^{-2/3}\right).$
@@
$\sum_{t=E\cdot|\mathcal{A}|+1}^{T}\mathbbm{1}[t\not\in\cup_{a\in\mathcal{A}_{\epsilon^{*}}}S(a)]=\sum_{a\not\in\mathcal{A}_{\epsilon}}n_{E\cdot|\mathcal{A}|+1,T}(a)\leq\Theta\left((\log T)^{1/3}T^{2/3}|\mathcal{A}|^{1/3}|\mathcal{B}|^{-2/3}\right)$
@@
Appendix E Proofs for Section 5
E.1 Alignment and continuity discussion
@@
E.2 Proofs and examples for Section 5.1
@@
Let $\hat{v}_{1,t}(a)$ be the empirical mean specified in LipschitzUCB at the beginning of time step $t$, which is the mean of the leader’s stochastic rewards $\left\{r_{1,t^{\prime}}(a_{t^{\prime}},b_{t^{\prime}})\mid a_{t^{\prime}}=a,1\leq t^{\prime}<t\right\}$. We also define $\hat{v}_{1,t}(a,b)$ to be the mean of the leader’s stochastic rewards for the arm $(a,b)$ up through time step $t-1$ (the set given by $\left\{r_{1,t^{\prime}}(a_{t^{\prime}},b_{t^{\prime}})\mid a_{t^{\prime}}=a,b_{t^{\prime}}=b,1\leq t^{\prime}<t\right\}$). Note that this quantity is not computable by the leader since the algorithm operates in the strongly decentralized setting, but we nonetheless find it convenient to consider in the analysis. Let $n_{t}(a)=\left|1\leq t^{\prime}<t\mid a_{t}=a\right|$ be the number of times that $a$ has been chosen prior to time step $t$. Let $n_{t}(a,b)=\left|1\leq t^{\prime}<t\mid a_{t}=a,b_{t}=b\right|$ be the number of times that $(a,b)$ has been chosen prior to time step $t$. For each arm $a\in\mathcal{A}$, let $b^{*}(a)=\text{argmax}_{b\in\mathcal{B}}v_{2}(a,b)$ be the follower’s best response.
@@
$\forall a\in\mathcal{A},t\leq T:\sum_{1\leq t^{\prime}<t\mid a_{t}=a}(v_{2}(a,b^{*}(a))-v_{2}(a_{t},b_{t}))\leq C^{\prime}\sqrt{n_{t}(a)\log T}$
@@
$\left|\hat{v}_{1,t}(a)-v_{1}(a,b^{*}(a))\right|\leq\frac{10\sqrt{\mathcal{B}\log T}}{\sqrt{n_{t}(a)}}+C^{\prime}\cdot L\cdot\frac{\sqrt{|\mathcal{B}|\log T}}{\sqrt{n_{t}(a)}}.$
@@
$B_{1}$ $:=\sum_{t=1}^{T}\left(\frac{10\sqrt{\mathcal{B}\log T}}{\sqrt{n_{t}(a)}}+C^{\prime}\cdot L\cdot\frac{\sqrt{|\mathcal{B}|\log T}}{\sqrt{n_{t}(a)}}\right)\leq O\left(L\sqrt{T|\mathcal{A}||\mathcal{B}|\log T}\right)$
@@
$=\sum_{a\in\mathcal{A}}\sum_{t\in T\mid a_{t}=a}\left(v_{1}(a,b^{*}(a))-v_{1}(a,b_{t})\right)$
@@
$=C^{\prime}\cdot\sqrt{\mathcal{B}|\log T}\cdot\sum_{a\in\mathcal{A}}\sqrt{n_{t}(a)}$
@@
$B_{1}$ $=\sum_{t=1}^{T}\left(\frac{10\sqrt{\mathcal{B}\log T}}{\sqrt{n_{t}(a)}}+C^{\prime}\cdot L\cdot\frac{\sqrt{|\mathcal{B}|\log T}}{\sqrt{n_{t}(a)}}\right)$
$=\sum_{t=1}^{T}\frac{10\sqrt{\mathcal{B}\log T}}{\sqrt{n_{t}(a)}}+\sum_{t=1}^{T}C^{\prime}\cdot L\cdot\frac{\sqrt{|\mathcal{B}|\log T}}{\sqrt{n_{t}(a)}}$
$\leq(10\sqrt{|\mathcal{B}\log T}+C^{\prime}\cdot L\sqrt{|\mathcal{B}|\log T})\sum_{t=1}^{T}\frac{1}{\sqrt{n_{t}(a)}}$
@@
where (A) follows from Lemma B.1
@@
Assume that clean event $G$ holds. This occurs with probability at least $1-(|\mathcal{A}+1)T^{-3}$ (Lemma E.1), so the clean event not occurring counts negligibly towards regret.
@@
$=2\cdot\sum_{t=1}^{T}\left(\frac{10\sqrt{\mathcal{B}\log T}}{\sqrt{n_{t}(a)}}+C^{\prime}\cdot L\cdot\frac{\sqrt{|\mathcal{B}|\log T}}{\sqrt{n_{t}(a)}}\right)+L\cdot B_{2}$
@@
where (A) uses Lemma E.2 and (B) uses Lemma E.3.
@@
$=2L\cdot\sum_{t=1}^{T}\left(\frac{10\sqrt{\mathcal{B}\log T}}{\sqrt{n_{t}(a)}}+C^{\prime}\cdot L\cdot\frac{\sqrt{|\mathcal{B}|\log T}}{\sqrt{n_{t}(a)}}\right)+B_{2}$
@@
where (A) uses the fact that $a^{*}$ is the action chosen by the leader at the Stackelberg equilibrium where (B) uses Lemma E.2 and (C) uses Lemma E.3.
E.3 Proof of Theorem 5.3
@@
for $(a,b)$ up to time step $t$. (The leader can observe this information because this algorithm operates in the weakly decentralized setting.) Let $v^{\text{UCB}}_{1,t}(a,b)$ denote the UCB estimate in PhasedUCB during time step $t$. Let $n_{t}(a)=|\left\{1\leq t^{\prime}<t\mid a_{t}=a\right\}|$ be the number of times that arm $a$ is pulled before time step $t$. Let $n_{t}(a,b)=|\left\{1\leq t^{\prime}<t\mid a_{t}=a,b_{t}=b\right\}|$ be the number of times that arms $(a,b)$ are pulled before time step $t$. Let $C$ be a constant such that ActiveArmElimination has high-probability instantaneous regret $g(t,T,\mathcal{B})=C\cdot\sqrt{|\mathcal{B}|\log T/t)}$ (such a constant $C$ exists by Proposition 7.1). Let $\mathcal{B}_{t}(a)$ be the computation of the active set at line 3 of PhasedUCB during time step $t$. Let $s_{t}(a)$ be the value of the variable $s^{\prime}(a)$ at the end of the ComputeActiveArms algorithm, when it is called at the beginning of time step $t$ in PhasedUCB. Let $(a^{*},b^{*})$ be the Stackelberg equilibrium.
@@
where (A) uses the event $G_{L,F}$, (B) uses the event $G_{L}$, (C) applies the lower bound in Lemma E.5, (D) uses the upper bound in Lemma E.5 to see that:

### v2

@@ Appendix C Proofs for Section 4
The fact that $E^{\prime}<E$ means that the leader’s exploration phase takes place entirely during the follower’s exploration phase. Moreover, since the leader’s exploration parameter $E^{\prime}\cdot|\mathcal{B}|$ is divisible by $|\mathcal{B}|$, for every arm $a\in\mathcal{A}$, the follower pulls every arm $b\in\mathcal{B}$ an equal number of times. Given that follower explores evenly between the two arms $b_{1}$ and $b_{2}$, the leader’s expected average reward $\mathbb{E}[\hat{v}_{1}(a_{1})]$ from $a_{1}$ during the first $E^{\prime}\cdot|\mathcal{B}|$ rounds is given by $(0.6+0.2)/2=0.4$ and the leader’s expected average reward $\mathbb{E}[\hat{v}_{1}(a_{2})]$ average reward from $a_{2}$ is given by $(0.5+0.4)/2=0.45$.
The proofs boils down to analyzing the relationship between the distributions $\hat{v}_{1}(a_{1})$ and $\hat{v}_{1}(a_{2})$. Note that we allow $E,E^{\prime}$ to be arbitrary, so we cannot use standard concentration bounds. Instead, we leverage the symmetry of the distribution of the empirical mean $\hat{v}_{1}(a_{1})$ (this follows from the fact that $\hat{v}_{1}(a_{1})-\mathbb{E}[\hat{v}_{1}(a_{1})]$ is distributed as a Gaussian). This means that:
@@
Assume that the clean event $G$ holds. This occurs with probability at least $1-(|\mathcal{A}|\cdot|\mathcal{B}|+|\mathcal{A}|)T^{-3}$ (Lemma C.1), so the clean event not occuring counts negligibly towards regret.
@@
We focus on $t>E_{2}\cdot|\mathcal{B}|\cdot|\mathcal{A}|+E_{1}\cdot|\mathcal{A}|$ for the remainder of the analysis. Our main ingredient is Lemma C.2. Note that $\epsilon^{*}=\Theta\left(\max\left(\frac{\sqrt{\log T}}{\sqrt{E_{1}}},\frac{\sqrt{\log T}}{\sqrt{E_{2}}}\right)\right)=\Theta\left((|\mathcal{A}|\cdot|\mathcal{B}|\cdot(\log T))^{1/3}\cdot T^{-1/3}\right)$ based on the settings of $E_{1}$ and $E_{2}$. The regret of the leader can be bounded as:
@@
where (A) follows from Lemma C.2. The regret of the follower can similarly be bounded as:
@@
where (B) follows from Lemma C.2. This proves the desired result.
C.3 Proof of Theorem 4.5
@@
Lemma C.3.
Assume the notation above. Let $\texttt{ALG}_{2}$ be any algorithm with high-probability instantaneous regret $g$ where $g(t,T,\mathcal{B})=O(E^{-1/2}|\mathcal{B}|^{1/2}(\log T)^{1/2})$ for $t>E$ and $g(t,T,\mathcal{B})=1$ for $t\leq E$, and let $\texttt{ALG}_{1}=\texttt{ExploreThenUCB}(E)$. Then, the event $G$ occurs with high probability: $\mathbb{P}[G]\geq 1-T^{-3}(|\mathcal{A}|+1)$.
@@
Lemma C.4.
Assume the notation above. Suppose that the clean event $G$ holds. Then it holds that:
$\frac{1}{n_{E\cdot|\mathcal{A}|,T+1}(a^{\prime})-1}\sum_{t\in S(a^{\prime})\setminus\left\{\max(S(a^{\prime}))\right\}}v_{1}(a_{t},b_{t})\geq\max_{a\in\mathcal{A}}\min_{b\in\mathcal{B}_{\epsilon^{*}}(a)}v_{1}(a,b)-\frac{20\sqrt{\log T}}{\sqrt{n_{E\cdot|\mathcal{A}|,T+1}(a^{\prime})}-1}.$
@@
$\frac{1}{n_{E\cdot|\mathcal{A}|,T+1}(a^{\prime})-1}\sum_{t\in S(a^{\prime})\setminus\left\{t^{*}\right\}}v_{1}(a_{t},b_{t})$ $\geq_{(1)}\hat{v}_{1,t^{*}}(a^{\prime})-\frac{10\sqrt{\log T}}{\sqrt{n_{E\cdot|\mathcal{A}|,T+1}(a^{\prime})-1}}$
$\geq v^{\text{UCB}}_{1,t^{*}}(a^{\prime})-\frac{20\sqrt{\log T}}{\sqrt{n_{E\cdot|\mathcal{A}|,T+1}(a^{\prime})-1}}$
$\geq v^{\text{UCB}}_{1,t^{*}}(\tilde{a})-\frac{20\sqrt{\log T}}{\sqrt{n_{E\cdot|\mathcal{A}|,T+1}(a^{\prime})-1}}$
$=\hat{v}_{1,t^{*}}(\tilde{a})+\frac{10\sqrt{\log T}}{\sqrt{|S|}}-\frac{20\sqrt{\log T}}{\sqrt{n_{E\cdot|\mathcal{A}|,T+1}(a^{\prime})-1}}$
$\geq_{(2)}\frac{1}{|S|}\sum_{t\in S}v_{1}(a_{t},b_{t})-\frac{20\sqrt{\log T}}{\sqrt{n_{E\cdot|\mathcal{A}|,T+1}(a^{\prime})-1}}$
$\geq_{(3)}\max_{a\in\mathcal{A}}\min_{b\in\mathcal{B}_{\epsilon^{*}}(a)}v_{1}(a,b)-\frac{20\sqrt{\log T}}{\sqrt{n_{E\cdot|\mathcal{A}|,T+1}(a^{\prime})-1}}$
@@
Assume that the clean event $G$ holds. This occurs with probability at least $1-(1+|\mathcal{A}|)T^{-3}$ (Lemma C.3), so the clean event not occurring counts negligibly towards regret.
@@
$\beta_{1}^{\text{tol}}\cdot(T-E\cdot\left|\mathcal{A}\right|)-\sum_{t=E\cdot|\mathcal{A}|+1}^{T}v_{1}(a_{t},b_{t})$
@@
$=\sum_{a\in\mathcal{A}}\sum_{t\in S(a)}\left(\epsilon^{*}+\max_{a\in\mathcal{A}}\min_{b\in\mathcal{B}_{\epsilon^{*}}(a)}v_{1}(a,b)-v_{1}(a_{t},b_{t})\right)$
$\leq|\mathcal{A}|+\sum_{a\in\mathcal{A}}\sum_{t\in S(a)\setminus\left\{\max(S(a))\right\}}\left(\epsilon^{*}+\max_{a\in\mathcal{A}}\min_{b\in\mathcal{B}_{\epsilon^{*}}(a)}v_{1}(a,b)-v_{1}(a_{t},b_{t})\right)$
@@
$+\underbrace{\sum_{a\in\mathcal{A}}(n_{E\cdot|\mathcal{A}|,T+1}(a)-1)\cdot\left(\max_{a\in\mathcal{A}}\min_{b\in\mathcal{B}_{\epsilon^{*}}(a)}v_{1}(a,b)-\frac{1}{n_{E\cdot|\mathcal{A}|,T+1}(a)-1}\sum_{t\in S(a)\setminus\left\{\max(S(a))\right\}}\left(v_{1}(a_{t},b_{t})\right)\right)}_{(2)}$
@@
$\sum_{a\in\mathcal{A}}(n_{E\cdot|\mathcal{A}|,T+1}(a)-1)\cdot\left(\max_{a\in\mathcal{A}}\min_{b\in\mathcal{B}_{\epsilon^{*}}(a)}v_{1}(a,b)-\frac{1}{n_{E\cdot|\mathcal{A}|,T+1}(a)-1}\sum_{t\in S(a)\setminus\left\{\max(S(a))\right\}}\left(v_{1}(a_{t},b_{t})\right)\right)$
$\leq\sum_{a\in\mathcal{A}}(n_{E\cdot|\mathcal{A}|,T+1}(a)-1)\cdot\frac{20\sqrt{\log T}}{\sqrt{n_{E\cdot|\mathcal{A}|,T+1}(a)}-1}$
$\leq O\left(\sqrt{\log T}\cdot\sum_{a\in\mathcal{A}}\sqrt{n_{E\cdot|\mathcal{A}|,T+1}(a)-1}\right)$
@@
where the first inequality uses Lemma C.4 and the last inequality uses Jensen’s inequality.
@@
$\frac{1}{n_{E\cdot|\mathcal{A}|,T+1}(a^{\prime})-1}\sum_{t\in S(a^{\prime})\setminus\left\{\max(S(a^{\prime}))\right\}}v_{1}(a_{t^{\prime}},b_{t^{\prime}})<\max_{a\in\mathcal{A}}\min_{b\in\mathcal{B}_{\epsilon^{*}}(a)}v_{1}(a,b)-\epsilon^{*}.$
However, by Lemma C.4, we can also lower bound the average reward across all time steps (except for the last time step) where $a^{\prime}$ is pulled in terms of $n_{E\cdot|\mathcal{A}|,T+1}(a^{\prime})$ as follows:
$\frac{1}{n_{E\cdot|\mathcal{A}|,T+1}(a^{\prime})-1}\sum_{t\in S(a^{\prime})\setminus\left\{\max(S(a^{\prime}))\right\}}v_{1}(a_{t^{\prime}},b_{t^{\prime}})\geq\max_{a\in\mathcal{A}}\min_{b\in\mathcal{B}_{\epsilon^{*}}(a)}v_{1}(a,b)-\frac{10\sqrt{\log T}}{\sqrt{n_{E\cdot|\mathcal{A}|,T+1}(a^{\prime})-1}}.$
@@
$\frac{10\sqrt{\log T}}{\sqrt{n_{E\cdot|\mathcal{A}|,T+1}(a^{\prime})-1}}\geq\epsilon^{*},$
@@
$n_{E\cdot|\mathcal{A}|,T+1}(a^{\prime})\leq\Theta\left(\frac{\log T}{(\epsilon^{*})^{2}}\right)=\Theta\left((\log T)^{1/3}T^{2/3}|\mathcal{A}|^{-2/3}|\mathcal{B}|^{-2/3}\right).$
@@
$\sum_{t=E\cdot|\mathcal{A}|+1}^{T}\mathbbm{1}[t\not\in\cup_{a\in\mathcal{A}_{\epsilon^{*}}}S(a)]=\sum_{a\not\in\mathcal{A}_{\epsilon}}n_{E\cdot|\mathcal{A}|,T+1}(a)\leq\Theta\left((\log T)^{1/3}T^{2/3}|\mathcal{A}|^{1/3}|\mathcal{B}|^{-2/3}\right)$
@@
Appendix D Proofs for Section 5
D.1 Alignment and continuity discussion
@@
D.2 Proofs and examples for Section 5.1
@@
Let $\hat{v}_{1,t}(a)$ be the empirical mean specified in LipschitzUCB at the beginning of time step $t$, which is the mean of the leader’s stochastic rewards $\left\{r_{1,t^{\prime}}(a_{t^{\prime}},b_{t^{\prime}})\mid a_{t^{\prime}}=a,1\leq t^{\prime}<t\right\}$. We also define $\hat{v}_{1,t}(a,b)$ to be the mean of the leader’s stochastic rewards for the arm $(a,b)$ up through time step $t-1$ (the set given by $\left\{r_{1,t^{\prime}}(a_{t^{\prime}},b_{t^{\prime}})\mid a_{t^{\prime}}=a,b_{t^{\prime}}=b,1\leq t^{\prime}<t\right\}$). Note that this quantity is not computable by the leader in a StrongDSG, but we nonetheless find it convenient to consider in the analysis. Let $n_{t}(a)=\left|1\leq t^{\prime}<t\mid a_{t}=a\right|$ be the number of times that $a$ has been chosen prior to time step $t$. Let $n_{t}(a,b)=\left|1\leq t^{\prime}<t\mid a_{t}=a,b_{t}=b\right|$ be the number of times that $(a,b)$ has been chosen prior to time step $t$. For each arm $a\in\mathcal{A}$, let $b^{*}(a)=\text{argmax}_{b\in\mathcal{B}}v_{2}(a,b)$ be the follower’s best response.
@@
$\forall a\in\mathcal{A},t\leq T:\sum_{1\leq t^{\prime}<t\mid a_{t}=a}(v_{2}(a,b^{*}(a))-v_{2}(a_{t},b_{t}))\leq C^{\prime}\sqrt{|\mathcal{B}|n_{t}(a)\log T}$
@@
$\left|\hat{v}_{1,t}(a)-v_{1}(a,b^{*}(a))\right|\leq\frac{10\sqrt{\left|\mathcal{B}\right|\log T}}{\sqrt{n_{t}(a)}}+C^{\prime}\cdot L\cdot\frac{\sqrt{|\mathcal{B}|\log T}}{\sqrt{n_{t}(a)}}.$
@@
$B_{1}$ $:=\sum_{t=1}^{T}\left(\frac{10\sqrt{\mathcal{B}\log T}}{\sqrt{n_{t}(a_{t})}}+C^{\prime}\cdot L\cdot\frac{\sqrt{|\mathcal{B}|\log T}}{\sqrt{n_{t}(a_{t})}}\right)\leq O\left(L\sqrt{T|\mathcal{A}||\mathcal{B}|\log T}\right)$
@@
$=\sum_{a\in\mathcal{A}}\sum_{t\in T\mid a_{t}=a}\left(v_{2}(a,b^{*}(a))-v_{2}(a,b_{t})\right)$
@@
$=C^{\prime}\cdot\sqrt{|\mathcal{B}|\log T}\cdot\sum_{a\in\mathcal{A}}\sqrt{n_{T}(a)}$
@@
$B_{1}$ $=\sum_{t=1}^{T}\left(\frac{10\sqrt{|\mathcal{B}|\log T}}{\sqrt{n_{t}(a_{t})}}+C^{\prime}\cdot L\cdot\frac{\sqrt{|\mathcal{B}|\log T}}{\sqrt{n_{t}(a_{t})}}\right)$
$=\sum_{t=1}^{T}\frac{10\sqrt{|\mathcal{B}|\log T}}{\sqrt{n_{t}(a_{t})}}+\sum_{t=1}^{T}C^{\prime}\cdot L\cdot\frac{\sqrt{|\mathcal{B}|\log T}}{\sqrt{n_{t}(a_{t})}}$
$\leq(10\sqrt{|\mathcal{B}|\log T}+C^{\prime}\cdot L\sqrt{|\mathcal{B}|\log T})\sum_{t=1}^{T}\frac{1}{\sqrt{n_{t}(a_{t})}}$
@@
where (A) follows from Lemma A.1
@@
Assume that clean event $G$ holds. This occurs with probability at least $1-(|\mathcal{A}+1)T^{-3}$ (Lemma D.1), so the clean event not occurring counts negligibly towards regret.
@@
$=2\cdot\sum_{t=1}^{T}\left(\frac{10\sqrt{\mathcal{B}\log T}}{\sqrt{n_{t}(a_{t})}}+C^{\prime}\cdot L\cdot\frac{\sqrt{|\mathcal{B}|\log T}}{\sqrt{n_{t}(a_{t})}}\right)+L\cdot B_{2}$
@@
where (A) uses Lemma D.2 and (B) uses Lemma D.3.
@@
$=2L\cdot\sum_{t=1}^{T}\left(\frac{10\sqrt{\mathcal{B}\log T}}{\sqrt{n_{t}(a_{t})}}+C^{\prime}\cdot L\cdot\frac{\sqrt{|\mathcal{B}|\log T}}{\sqrt{n_{t}(a_{t})}}\right)+B_{2}$
@@
where (A) uses the fact that $a^{*}$ is the action chosen by the leader at the Stackelberg equilibrium where (B) uses Lemma D.2 and (C) uses Lemma D.3.
D.3 Proof of Theorem 5.3
@@
for $(a,b)$ up to time step $t$. (The leader can observe this information in a WeakDSG.) Let $v^{\text{UCB}}_{1,t}(a,b)$ denote the UCB estimate in PhasedUCB during time step $t$. Let $n_{t}(a)=|\left\{1\leq t^{\prime}<t\mid a_{t}=a\right\}|$ be the number of times that arm $a$ is pulled before time step $t$. Let $n_{t}(a,b)=|\left\{1\leq t^{\prime}<t\mid a_{t}=a,b_{t}=b\right\}|$ be the number of times that arms $(a,b)$ are pulled before time step $t$. Let $C$ be a constant such that ActiveArmElimination has high-probability instantaneous regret $g(t,T,\mathcal{B})=C\cdot\sqrt{|\mathcal{B}|\log T/t)}$ (such a constant $C$ exists by Proposition 7.1). Let $\mathcal{B}_{t}(a)$ be the computation of the active set at line 3 of PhasedUCB during time step $t$. Let $s_{t}(a)$ be the value of the variable $s^{\prime}(a)$ at the end of the ComputeActiveArms algorithm, when it is called at the beginning of time step $t$ in PhasedUCB. Let $(a^{*},b^{*})$ be the Stackelberg equilibrium.
@@
where (A) uses the event $G_{L,F}$, (B) uses the event $G_{L}$, (C) applies the lower bound in Lemma D.5, (D) uses the upper bound in Lemma D.5 to see that: